teressa24:

Variable x is 7 more than variable y. Variable x is also 1 less than y. Which of the following pairs of equations best models the relationship between x and y? (4 points) x = 1y x = y + 7 x = y + 7 x = y − 1 y = x + 8 y = x − 1 y = 7x y = x + 1

jhonyy9:

any idea ? what mean x is 7 more than y ?

jhonyy9:

or x is 1 less than y ? how you write these ?

jhonyy9:

@Laylalyssa

teressa24:

idk

TypicalFoxGirl:

If x is 7 more than y and then x is 1 less than y Then the obvious answer is: x = y + 7 x = y − 1 as it has both statements.

TypicalFoxGirl:

@jhonyy9 if you were having trouble, there you go-